OpenStudy (anonymous):

Find the discriminant of the quadratic equation below. x^2 + 4x + 9 = 0

OpenStudy (anonymous):

For general quadratic equation, Ax^2 +Bx+C discriminant is given as \[\sqrt{B^{2}-4*A*C}\]

OpenStudy (anonymous):

\[1x^2+4x+9=0\] \[A=1, B=4, C=9\] Use the equation above to find the discriminant

OpenStudy (anonymous):

im gonna ask a stupid question but what is a discriminant?

OpenStudy (anonymous):

It aint a stupid question The discriminant is an equation that tells us how many real roots our polynomial has So if the discriminant is greater than 0 then our polynomial has 2 real roots If it equals 0 then it has only one real root If it equals less than 0 then there are no real roots

OpenStudy (anonymous):

The equation of the discriminant is : \[\sqrt{B^{2}-4*A*C}\] I forgot to mention that a polynomial is an equation with the highest degree being 2: \[y=3x^{\color{orange}2}+5x+3\] Now a polynomial in standard form looks as follows: \[Ax^2+Bx+C=0\] where A, B and C are numerical values So in our case the equation is \[\color{orange}{1}x^2+\color{orange}{4}x+\color{orange}{9}=0\] So our A=1, B=4 and C=9

OpenStudy (anonymous):

Now we need to plug this info into our discriminant equation \[Discriminant= \sqrt{4^2-4*(1)(9)}\] \[=\sqrt{16-36}\] \[=-20\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Its kinda simple once you memorize the equation

OpenStudy (anonymous):

yeah it sure is

OpenStudy (anonymous):

LOL GL